OpenStudy (anonymous):

A radioactive isotope (239)Pu has a half life of 24,360 years. What is the initial amount of the isotope if there is only 4.2 grams left after 1,000 years? please show all work\

OpenStudy (anonymous):

Y(t)= = Y(0) e^kt Y(0) = ?

OpenStudy (anonymous):

sub the values for the time vs the size

OpenStudy (anonymous):

so Y(t)=4.2 when t =1000

OpenStudy (anonymous):

substitute the value of the half life at t=24360

OpenStudy (anonymous):

\[4.2=Q_0\times (\frac{1}{2})^{\frac{1000}{24360}}\] \[Q_0=4.2\div(\frac{1}{2})^{\frac{1000}{24360}}\]